### v1

# One Step is Enough: Multi-Agent Reinforcement Learning based on One-Step Policy Optimization for Order Dispatch on Ride-Sharing Platforms

Abstract
On-demand ride-sharing platforms face the fundamental challenge of dynamically bundling passengers with diverse origins and destinations and matching them with vehicles in real time, all under significant uncertainty. Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution for this problem, leveraging decentralized learning to address the curse of dimensionality caused by the large number of agents in the ride-hailing market and the resulting expansive state and action spaces. However, conventional MARL-based ride-sharing approaches heavily rely on the accurate estimation of Q-values or V-values, which becomes problematic in large-scale, highly uncertain environments. Specifically, most of these approaches adopt an independent paradigm, exacerbating this issue, as each agent treats others as part of the environment, leading to unstable training and substantial estimation bias in value functions. To address these challenges, we propose two novel alternative methods that bypass value function estimation. First, we adapt Group Relative Policy Optimization (GRPO) to ride-sharing, replacing the Proximal Policy Optimization (PPO) baseline with the group average reward to eliminate critic estimation errors and reduce training bias. Second, inspired by GRPO’s full utilization of group reward information, we customize the PPO framework for ride-sharing platforms and show that, under a homogeneous fleet, the optimal policy can be trained using only one-step rewards—a method we term One-Step Policy Optimization (OSPO). Experiments on a real-world Manhattan ride-hailing dataset demonstrate that both GRPO and OSPO achieve superior performance across most scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks. Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded simulation time of GRPO. Our code, trained models, and processed data are publicly available at the anonymous repository: https://github.com/RS2002/OSPO.
Introduction
The widespread adoption of on-demand ride-sharing platforms has fundamentally transformed urban transportation systems, offering scalable and sustainable mobility solutions for modern cities. By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (Wang and Yang 2019; Santi et al. 2014). However, managing such platforms presents significant operational complexity. Passengers arrive randomly over time, each with unique origins and destinations, and the platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suitable vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms.
Recently, reinforcement learning (RL) methods have shown great promise in addressing the order dispatch problem on ride-sharing platforms. Although ride-sharing is fundamentally a centralized task—requiring the platform to aggregate information from all drivers and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (Sivagnanam et al. 2024)) pose a serious curse of dimensionality, rendering single-agent RL approaches intractable. As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each driver is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives. MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (Jin et al. 2025). Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to single-agent RL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents. However, DTDE methods rely on the independent assumption that each agent acts based only on its local information and treats other agents as part of the environment. This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (Hu, Feng, and Li 2025), ultimately resulting in suboptimal and inefficient learning.
To address this problem, we propose two multi-agent policy optimization methods that do not require estimating Q-values or V-values. First, we adapt Group Relative Policy Optimization (GRPO) (Shao et al. 2024) for the ride-sharing task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (Schulman et al. 2017) with the group average reward. This modification significantly reduces the computational resources required for training and minimizes bias by eliminating neural network estimation errors associated with V-values and Q-values. Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step rewards by leveraging the homogeneous characteristics of agents in the ride-sharing task. Both methods are validated through case studies using real-world ride-hailing data and demonstrate superior performance compared to state-of-the-art approaches. The main contributions of this paper are summarized as follows:
  • We introduce a novel ride-sharing framework based on GRPO. By replacing the baseline in PPO with group average reward, our method eliminates the need to estimate V-values. It successfully improves training performance and stability by removing the influence of estimation bias. To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing platforms.
  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in ride sharing task, the advantage function can be replaced by the group relative reward within only one step. To the best of our knowledge, it is the simplest and most efficient RL method for the ride-sharing task. Furthermore, we also specify the conditions under which OSPO can be applied in identical MARL scenarios.
  • We validate the proposed methods using a real-world ride-hailing dataset from Manhattan. Experimental results show that both GRPO and OSPO consistently outperform existing ride-sharing approaches as well as conventional Policy Gradient (PG) methods across most scenarios, primarily by optimizing pickup times and increasing the number of served orders. Notably, both methods maintain high efficiency even when implemented with a simple Multilayer Perceptron (MLP) network, with OSPO achieving the lowest GPU utilization among all methods. Furthermore, OSPO surpasses GRPO in performance, which can be attributed to its elimination of the bias introduced by the bounded simulation time present in GRPO.
Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each driver-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using PPO-style policy gradients.
Preliminary
Group Relative Policy Optimization (GRPO)
With the success and popularity of Reinforcement Learning from Human Feedback (RLHF), RL-based LLM post-training has garnered increasing attention. This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (Rafailov et al. 2023), Reinforce Leave-One-Out (RLOO) (Ahmadian et al. 2024), and GRPO (Shao et al. 2024). Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (Shao et al. 2024) and DeepSeek-V3 (Liu et al. 2024).
Serving as a PG (Sutton et al. 1999) based method, the optimization objective of GRPO has a similar format to PPO(Schulman et al. 2017), expressed as:
$\text{J}_{GRPO}(\theta)=\mathbb{E}_{s,u\sim\pi_{\theta^{-}}}\left[\min\left\{\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)}\hat{\text{A}}_{\pi_{\theta}}(s,u),\text{CLIP}\right.\right.$ (1)
$\left.\left(\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg{]}-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (Yu et al. 2025), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
Aside from the KL divergence term, the core difference between GRPO and PPO lies in the definition of the advantage function. In PPO, the original advantage function is defined as:
$\text{A}_{\pi}(s_{t},u_{t})=\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t})\ ,$ (2)
where the Q-value can be estimated using real trajectories, and the V-value can be estimated through a critic network. In contrast, GRPO proposes a simpler and more efficient advantage function given by:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\sum_{\tau\in K_{t}}\gamma^{\tau}\frac{\text{R}(s_{t+\tau},u_{t+\tau})-\mu}{\sigma}\ ,$ (3)
[…]
Problem Setup for Ride-sharing Order Dispatch
In this paper, we consider the optimal order assignment task of an on-demand ride-sharing platform with $n$ homogeneous drivers (agents), each equipped with a vehicle of capacity $c$. At each time step $t$, there will be $m_{t}$ newly arrived orders from customers in different locations. The platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suit- able vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (Littman 1994), defined as $<n,S,U,P,R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor. The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e. information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$. In our ride-sharing task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and remaining delivery time. The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$. At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$. The transition probability function $P$ is not modeled explicitly in model-free RL methods. For the reward function $\text{r}(s_{i},u_{i})$, we leave the detailed design in appendix. Since ride-sharing is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
Methodology
In this section, we first describe how we adapt GRPO for the multi-agent ride-sharing scenario, and analyze the theoretical foundations that support its applicability. We then introduce our proposed OSPO method and specify the conditions under which it can be effectively applied.
GRPO for Ride Sharing
The workflow of GRPO for ride sharing is shown in Fig. 1, following a design similar to previous SAC-based work (Enders et al. 2023). Firstly, we require the policy network $\pi_{\theta}$ to directly output the action probability of each agent, expressed as:
$p_{i,t}=\pi_{\theta}([s_{i,t},s^{o}_{t}])\ ,$ (4)
where $s_{i,t}$ and $s^{o}_{t}$ represent the state of agent $i$ and the global state at time $t$, respectively, and $p_{i,t}\in[0,1]^{w_{t}}$ represents the action probability (i.e., the probability of matching each order to agent $i$). Due to the homogeneity of the agents, we utilize a shared network among all agents, similar to approaches taken in previous works.
Considering the action constraint that an order cannot be assigned to multiple agents, we cannot simply choose the action independently for each agent. Instead, we view the action probabilities as scores and use bipartite matching to maximize the assignment score, following (Enders et al. 2023; Hu, Feng, and Li 2025):
$\max_{U_{t}}\sum_{i\in\mathcal{I}}u_{i,j,t}\cdot y_{i,j,t},$ (5a)
s.t. $\sum_{i\in\mathcal{I}}u_{i,j,t}\leq 1,\quad\forall j\in\mathcal{J}_{t},$ (5b)
$\sum_{j\in\mathcal{J}_{t}}u_{i,j,t}\leq 1,\quad\forall i\in\mathcal{I},$ (5c)
$u_{i,j,t}\in\{0,1\},\quad\forall i\in\mathcal{I},j\in\mathcal{J}_{t},$ (5d)
where $u_{i,j,t}$ indicates whether agent $i$ is assigned order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=-\infty$ for all unavailable workers at time $t$). The set =$\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned. Constraint (5b) ensures that an order can be assigned to at most one agent, while constraint (5c) guarantees that each agent is assigned at most one order. After that, we can obtain the reward according to Eq. 13 and then optimize the policy by maximizing Eq. 1.
To make GRPO work in our homogeneous cooperative MARL scenario, only two minor changes should be made to the original objective function in the SARL scenario. First, when calculating the advantage function, we use the rewards of all agents over all steps to compute the mean value $\mu$ and standard deviation $\sigma$ in Eq. 3. (For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$. This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.) Second, for the KL normalization term in Eq. 1, since we do not have a reference policy like the LLM post-training, we propose to regularly track policy performance and retain the best checkpoint. This best policy serves as the reference policy, helping to ensure that the current policy does not deviate significantly from the best one. A detailed algorithm is provided in appendix.
Next, we analyze why our modification to GRPO is valid, allowing the transfer from SARL to MARL. For the advantage function (Eq. 3) in GRPO, we can express it as:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$ (6)
For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (Schulman et al. 2017). For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 00 as the number of sampled trajectories increases (i.e., as the group becomes larger).
The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (Wang et al. 2023). An important feature of such a reward function is that each step reward has a similar range, leading to: $\mathbb{E}_{s_{t},u_{t}\sim\pi}[\text{R}(s_{t},u_{t})]=\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi_{t}$, where $\xi_{t}$ is a small residual and $\mathbb{E}_{t}[\xi_{t}]=0$. Thus, we have:
$\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (7)
$=$ $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s_{\tau},u_{\tau}\sim\pi}[\text{R}(s_{\tau},u_{\tau})]-\xi_{\tau}+\xi)$
$=$ $\text{V}_{\pi}(s_{t})+\sum_{\tau\in K_{t}}\gamma^{\tau}(\xi-\xi_{\tau})\ .$
If we ignore the residual term, we obtain $\hat{\text{A}}_{\pi}(s_{t},u_{t})\approx\frac{1}{\sigma}(\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t}))$. Compared to the advantage in PPO (Eq. 2), there is only one additional factor $\frac{1}{\sigma}$. According to (Kurin et al. 2022), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training. Furthermore, based on the theory of policy gradients (Sutton et al. 1999), adding any baseline or positive factor does not influence the policy direction, regardless of the action. As a result, we consider the advantage of GRPO (Eq. 3) to be a valid replacement for the advantage in PPO (Eq. 2).
One Step Policy Optimization (OSPO)
We note that in the ride-sharability platform, all vehicles have identical capacity and are assigned to customers purely based on spatial proximity, without any additional attributes or preferences influencing the dispatch decisions. As a result, in the long run, the cumulative contribution or average earnings of each driver per unit time must converge to the same value. This is because the inherent randomness in order arrivals and assignments is experienced equally by all drivers over time. Therefore, it is reasonable to expect that different agents will exhibit similar value functions at each time step, regardless of their individual states. In this case, we have the following property:
$\text{V}_{\pi}(s_{1,t})+\epsilon_{1,t}=\text{V}_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (8)
$=\text{V}_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[\text{V}_{\pi}(s_{i,t})]\ ,$
where $\epsilon_{i,t}$ is a small residual term, with $\mathbb{E}_{i\in\mathcal{I}}[\epsilon_{i,t}]=0$. As a result, when we ignore the residual terms, the different agents do not differ significantly from the group in GRPO, making it feasible in our scenario. Based on the property in Eq. 8, we can recalculate the advantage function as follows:
$\text{A}_{\pi}(s_{i,t},u_{i,t})=\text{Q}_{\pi}(s_{i,t},u_{i,t})-\text{V}_{\pi}(s_{i,t})$ (9)
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\text{V}_{\pi}(s_{i,t+1})-\text{V}_{\pi}(s_{i,t})$
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\mathbb{E}_{j\in\mathcal{I},s_{j,t+1}\sim\pi}[\text{V}_{\pi}(s_{j,t+1})]$
$-\mathbb{E}_{j\in\mathcal{I},s_{j,t}\sim\pi}[\text{V}_{\pi}(s_{j,t})]+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}$
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\mathbb{E}_{j\in\mathcal{J},s_{j,t+1}\sim\pi}[\text{V}_{\pi}(s_{j,t+1})]$
$-\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t},s_{j,t+1}\sim\pi}[\text{r}(s_{j,t},u_{j,t})+\gamma\text{V}_{\pi}(s_{j+1,t})]$
$+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}$
$=$ $\text{r}(s_{i,t},u_{i,t})-\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}\ .$
If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (Kurin et al. 2022), and ignoring the residual term, we arrive at the advantage function:
$\hat{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu}{\sigma}\ .$ (10)
This indicates that we can use only the one-step reward for policy optimization, which we term One Step Policy Optimization (OSPO). Even though the format may seem counter-intuitive, we provide an intuitive explanation. The validity of Eq. 10 is based on Eq. 8: the V-values remain similar among agents at each step due to the homogeneity in the ride-sharing system. For any arbitrary agent, regardless of its actions in the current step, it can achieve a similar V-value as others in the next step. Therefore, under the independent assumption, the goal of each agent is to maximize its single-step reward. In Eq. 8, removing the baseline and factor leaves only the single-step reward $\text{r}(s_{i,t},u_{i,t})$, which aligns with our intuition. Compared to GRPO, our OSPO further improves computational efficiency. Moreover, our method demonstrates greater robustness since it is not influenced by the discount factor.
Additionally, we propose two changes to the advantage function. First, we suggest directly using the mean value and standard deviation at each step $t$ to replace the values over the entire episode:
$\bar{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (11)
This approach not only makes $\mu_{t}$ closer to $\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]$ in Eq. 9, but also further improves computational efficiency, achieving true one-step computation. Second, during training, to better align with the property described in Eq. 8, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$. This penalty is designed to reduce the utility gap (reflecting the V-value) between agents. Consequently, the final advantage function is expressed as:
$\tilde{\text{A}}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (12)
where $\alpha$ is a hyper-parameter. (Note that when Eq. 8 is satisfied, $\Delta\delta_{t}$ is close to 0. We also include the term $-\alpha\Delta\delta_{t}$ in the advantage function of GRPO.)
Experiment
Experiment Setup
To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (Taxi and Commission 2024). The data collected from 19:00 to 19:30 on July 17, 2024, is used as the training set, which includes 3,726 valid orders. The data from other periods of that day is used as the testing set.
For the simulation, we set the total number of drivers to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h. The optimal routing is achieved using the OSRM simulator (Luxen and Vetter 2011). The experiments are conducted using the PyTorch framework (Paszke et al. 2019) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. Due to device limitations, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode. We maximize the training episodes at 1,000 and evaluate model performance every 10 episodes.
To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (Maas et al. 2013) as the activation function. Considering the variable action space (since the number of orders changes each time), we set the network input as a single driver-order pair, ignoring the relationships between orders. Finally, a softmax layer is used to normalize the output probabilities.
During training, our models utilized approximately 3.8GB (OSPO) and 5.2GB (GRPO) of GPU memory, with a batch size of 256. For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (Hu, Feng, and Li 2025). For optimization, we use the Adam optimizer (Kingma 2014) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
In the following, we first compare our GRPO and OSPO ride-sharing methods with previously popular methods of different types, including DTDE, CTDE, and CTCE. Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (De Witt et al. 2020), Multi-Agent PPO (MAPPO) (Yu et al. 2022), and Independent PG (IPG) (Sutton et al. 1999).
Comparative Experiment
Figure 2: Training Process: Each method is trained three times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the range of fluctuations, while the solid line indicates the average value. (For delivery time and detour time, only completed orders are counted, as these metrics are uncertain for unfinished orders.)
Table 1: Model Performance on Testing Sets: Bold entries represent the best results.
Method Configuration Order Amount in Testing Sets Method Network GPU (GB) 2,850 3,114 3,577 3,910 DeepPool MLP (20K) 3.97 10,690.95±86.93 12,368.83±117.75 12,569.92±77.15 12,990.07±25.07 BMG-Q GAT (117K) 4.28 10,982.45±187.60 12,637.12±116.89 13,064.66±142.65 13,600.33±92.30 HIVES GRU (16M) 6.01 11,044.05±8.30 12,787.39±40.97 12,576.75±46.25 12,097.87±72.32 Enders et al. MLP (118K) 8.19 10,036.89±88.64 12,160.56±39.53 12,296.22±188.86 12,128.95±123.22 CEVD MLP (23K) 21.45 10,842.98±31.36 12,702.10±37.39 13,263.08±136.67 13,609.91±17.34 GRPO MLP (20K) 5.17 11,372.81±31.73 13,015.88±78.18 13,644.95±116.94 12,964.86±99.84 OSPO MLP (20K) 3.82 11,467.87±5.39 13,046.27±15.29 13,771.56±81.43 13,429.67±69.79
To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories. A more detailed overview of previous work is provided in the appendix.
  • DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019) (DTDE): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (Tan 1993; Van Hasselt, Guez, and Silver 2016) and using a CNN as the Q-network. To align it with our experimental scenario, we replace the CNN with an MLP similar to our methods (without the softmax).
  • BMG-Q (Hu, Feng, and Li 2025) (DTDE): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (Velickovic et al. 2017) to capture the relationships among neighboring agents.
  • HIVES (Hao and Varakantham 2022) (CTDE): Based on QMIX (Rashid et al. 2020), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
  • Enders et al. (Enders et al. 2023) (CTDE): Based on Multi-Agent Soft Actor Critc (MASAC) (Haarnoja et al. 2018), Enders et al. propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
  • CEVD (Bose et al. 2023) (CTCE): CEVD modifies the Value Decomposition (VD) paradigm (Sunehag et al. 2018) from CTDE to CTCE. Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
For fairness, we modify the state space and reward function to align closely with those used in this paper. Additionally, we also choose a Greedy method as a baseline, which directly matches the order to the closest driver. The training process and testing results are shown in Fig. 2 and Table 1, respectively. In Fig. 2, we illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order. We observe that the GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, allowing them to serve more orders. In Table 1, it is noteworthy that the OSPO, which has the smallest network size and lowest GPU utilization, achieves the best performance across three testing scenarios. However, we also find that the GRPO and OSPO do not perform as well when the order amount increases to 3,910, which is the only scenario where the order amount exceeds 3,726 in the training set. Notably, the two methods that perform relatively well in this scenario are BMG-Q (Hu, Feng, and Li 2025) and CEVD (Bose et al. 2023). The primary distinction of these methods compared to others is their use of neighbor information when calculating Q-values. In Fig. 2, we also observe that they achieve lower detour times, suggesting better cooperation among agents through neighbor information. These results imply that incorporating more global or neighbor information may be more beneficial in on-pick scenarios. Moving forward, it would be advantageous to explore the model’s performance when the training set includes multiple order amount scenarios.
Ablation Study
In the ablation study, we primarily investigate the effects of:
  • Reward Normalization: We compare the effect of using single-step reward normalization ($\mu_{t},\sigma_{t}$ in Eq. 11) versus using the whole episode reward ($\mu,\sigma$ in Eq. 10) to study the trade-off between computational efficiency and performance.
  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 12) to determine whether it helps to maintain property (ii) (Eq. 8).
  • KL Regulation Term: We assess the effect of using a KL regulation term ($\text{KL}(\pi_{\theta}\|\pi_{ref})$ in Eq. 1) to illustrate whether the historical best policy can prevent model divergence.
Figure 3: Result of Ablation Study
Since it seems unreasonable to use single-step reward normalization in GRPO, we omit this experiment. Additionally, we compare our methods with conventional policy optimization methods, including IPPO, MAPPO, and IPG. As shown in Fig. 3, we notice that the proposed GRPO and OSPO outperform all variant versions, illustrating the efficiency of the proposed modules. For the policy optimization methods, we observed that only IPPO demonstrates relatively good performance (even when compared to our proposed methods), while IPG struggles with high derivation and low sample efficiency. Additionally, MAPPO suffers from the curse of dimensionality in the centralized critic network, resulting in failure to converge.
Additionally, we observe that OSPO demonstrates overall better performance than GRPO. This is because, although the V-values among agents are generally similar, they have been maintained over a longer time period. In our ride-sharing scenario, the number of served orders may differ slightly among agents at each time step, but this variation can be considered negligible over the long term. Since our simulation lasts only 30 minutes, this difference becomes more pronounced when calculating the advantage (i.e., using the group average cumulative reward as a baseline introduces a slight bias). In contrast, OSPO exhibits greater robustness, as it relies solely on one-step rewards and does not encounter this issue.
Discussions
In this section, we analyze the current limitations and future directions of our work.
  • Application Constraints: As mentioned earlier, our method primarily relies on the property that the V-values of different agents remain similar at any time step (Eq. 8). This limits the application scenarios of OSPO to cooperative homogeneous environments. It cannot be used in cooperative scenarios where some agents must sacrifice their own interests to maximize the global reward. However, this condition can mostly be met in independent MARL, as each agent’s learning goal is to maximize its own long-term reward. In the next step, it is worthwhile to explore a wider range of possible application scenarios for OSPO.
  • Independent Assumption: One important reason for proposing OSPO is to reduce the estimation error of Q-values or V-values in independent MARL scenarios. However, the current training approach of our method still follows independent learning. The challenges of poor cooperation in independent MARL are yet to be addressed. Future work could explore combining OSPO with other methods, such as communication mechanisms or more powerful networks, similar to the approach taken by BMG-Q.
  • Fairness: Since OSPO is fundamentally based on the similar V-value property (Eq. 8), and we incorporate a punishment term (Eq. 12) to encourage the model to adhere to this property, it is important to explore whether this approach contributes to fairness. Recently, assignment fairness and discriminatory scenarios have garnered increasing attention in gig systems (Shi et al. 2021).
Conclusion
In this paper, we propose two novel MARL methods for ride-sharing that do not rely on estimating V-values or Q-values. First, we successfully adapt GRPO to the ride-sharing context, achieving promising performance by eliminating the estimation errors associated with the V-value network in PPO. Building on the insights from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards. To the best of our knowledge, OSPO is the simplest and most efficient reinforcement learning method for order dispatch on ride-sharing platforms to date. Extensive experiments on real-world ride-sharing data demonstrate that our methods consistently outperform previous approaches in most scenarios, while requiring only a simple MLP network and exhibiting low GPU utilization.
Appendix
Reward Function Design
The reward function is shaped following previous work (Hu, Feng, and Li 2025):
$\text{r}(s_{i},u_{i})$ =$=\begin{cases}\beta_{1}+\beta_{2}p^{in}_{i}-\beta_{3}p^{out}_{i}-\beta_{4}\chi_{i}-\beta_{5}\rho_{i},|u_{i}|=1\\ 0,|u_{i}|=0\end{cases}$ (13)
where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the payout to drivers, respectively. The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively. This reward function is designed to comprehensively consider the interests of the platform, drivers, and passengers, mimicking the operation of a real-world ride-sharing platform. It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively. When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route.
Algorithm
The detailed algorithm of our GRPO/OSPO for ride sharing method is illustrated in Algorithm 1.
Algorithm 1 GRPO for Ride Sharing
0:  Policy network $\pi_{\theta}$ with shared parameters $\theta$
0:  Best policy checkpoint $\pi_{\phi}$ (initialized as $\pi_{\theta}$)
0:  Environment simulator, Reward function (Eq. 13)
1:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
2:  for episode $=1$ to $M$ do
3:     Reset environment, get initial states $\{s_{i,0}\}_{i=1}^{n}$, $s^{o}_{0}$
4:     for time step $t=0$ to $T$ do
5:        for each agent $i\in\mathcal{I}$ do
6:           Compute action probabilities (Eq. 4)
7:        end for
8:        Solve assignment problem via BiPartite matching (Eq. 5), getting action $U_{t}$
9:        Execute assignment $U_{t}$, observe rewards $\{r_{i,t+1}\}_{i=1}^{n}$
10:        Store transition ($(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
11:     end for
12:     Policy Optimization:
13:     for epoch $=1$ to $M$ do
14:        Sample batch $\mathcal{B}\sim\mathcal{D}$
15:        Compute group advantage $\hat{A}_{t}$
16:        Update policy parameters $\theta$ by maximizing Eq. 1
17:     end for
18:     if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
19:        Update best checkpoint: $\phi\leftarrow\theta$
20:     end if
21:  end for
Related Work
In this section, we provide a brief overview of current ride-sharing methods. Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019), independent DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods. Subsequent works have primarily aimed to enhance algorithms by fostering better cooperation among agents. For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (Hu, Feng, and Li 2025). Li et al. (Li et al. 2019) also propose using mean field approaches to capture more useful global information.
Additionally, some works attempt to transfer the successes of CTDE and CTCE to ride-sharing, where centralized critics can facilitate better cooperation among agents. However, these approaches face various shortcomings and challenges. Lima et al. (De Lima et al. 2020) first introduce QMIX, but only consider tasks involving a limited number of drivers in simulated grid scenarios, which presents a significant gap to real-world applications. Hao et al. (Hao and Varakantham 2022) explore hierarchical structures to address the curse of dimensionality in QMIX (Rashid et al. 2020), but the input dimensions of the mixture network still increase with the number of agents. Bose et al. (Bose et al. 2023) propose a Value Decomposition (VD) (Sunehag et al. 2018) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (Zhang et al. 2024), order bundling (Jiang, Xu, and Varakantham 2025), and multi-modal transportation (Hu, Dong, and Li 2025). Since these topics extend beyond the scope of this paper, we do not discuss them further here.
Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (Hu, Feng, and Li 2025). Our work addresses this gap by proposing GRPO/OSPO—eliminating value estimation entirely through group rewards—for stable and scalable coordination in homogeneous MARL systems.

### v2

@@
 Abstract
-On-demand ride-sharing platforms face the fundamental challenge of dynamically bundling passengers with diverse origins and destinations and matching them with vehicles in real time, all under significant uncertainty.
-Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution for this problem, leveraging decentralized learning to address the curse of dimensionality caused by the large number of agents in the ride-hailing market and the resulting expansive state and action spaces.
-However, conventional MARL-based ride-sharing approaches heavily rely on the accurate estimation of Q-values or V-values, which becomes problematic in large-scale, highly uncertain environments.
-Specifically, most of these approaches adopt an independent paradigm, exacerbating this issue, as each agent treats others as part of the environment, leading to unstable training and substantial estimation bias in value functions.
-To address these challenges, we propose two novel alternative methods that bypass value function estimation.
-First, we adapt Group Relative Policy Optimization (GRPO) to ride-sharing, replacing the Proximal Policy Optimization (PPO) baseline with the group average reward to eliminate critic estimation errors and reduce training bias.
-Second, inspired by GRPO’s full utilization of group reward information, we customize the PPO framework for ride-sharing platforms and show that, under a homogeneous fleet, the optimal policy can be trained using only one-step rewards—a method we term One-Step Policy Optimization (OSPO).
-Experiments on a real-world Manhattan ride-hailing dataset demonstrate that both GRPO and OSPO achieve superior performance across most scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks.
-Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded simulation time of GRPO.
-Our code, trained models, and processed data are publicly available at the anonymous repository: https://github.com/RS2002/OSPO.
-Introduction
+Order dispatch is a critical task in ride-sharing systems with Autonomous Vehicles (AVs), directly influencing efficiency and profits.
+Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution to this problem by decomposing the large state and action spaces among individual agents, effectively addressing the Curse of Dimensionality (CoD) in transportation market, which is caused by the substantial number of vehicles, passengers, and orders.
+However, conventional MARL-based approaches heavily rely on accurate estimation of the value function, which becomes problematic in large-scale, highly uncertain environments.
+To address this issue, we propose two novel methods that bypass value function estimation, leveraging the homogeneous property of AV fleets.
+First, we draw an analogy between AV fleets and groups in Group Relative Policy Optimization (GRPO), adapting it to the order dispatch task.
+By replacing the Proximal Policy Optimization (PPO) baseline with the group average reward-to-go, GRPO eliminates critic estimation errors and reduces training bias.
+Inspired by this baseline replacement, we further propose One-Step Policy Optimization (OSPO), demonstrating that the optimal policy can be trained using only one-step group rewards under a homogeneous fleet.
+Experiments on a real-world ride-hailing dataset show that both GRPO and OSPO achieve promising performance across all scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks.
+Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded time horizon of GRPO.
+Our code, trained models, and processed data are provided at https://github.com/RS2002/OSPO.
+Machine Learning, ICML
+1 Introduction
 The widespread adoption of on-demand ride-sharing platforms has fundamentally transformed urban transportation systems, offering scalable and sustainable mobility solutions for modern cities.
-By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (Wang and Yang 2019; Santi et al.
-2014).
+By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (wang2019ridesourcing; santi2014quantifying).
+Recently, with the development of Autonomous Vehicle (AV) technology, systems like Robotaxi have been introduced by various companies, including Baidu, Didi, and Zoox.
+These systems not only offer lower operating costs through the AV fleet but also enable better management, as the fleet is fully controlled rather than relying on human drivers.
 However, managing such platforms presents significant operational complexity.
@@
 These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users.
-Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms.
+Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms with AVs.
 Recently, Reinforcement Learning (RL) methods have shown great promise in addressing the order dispatch problem on ride-sharing platforms.
-Although ride-sharing is fundamentally a centralized task—requiring the platform to aggregate information from all drivers and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (Sivagnanam et al.
-2024)) pose a serious curse of dimensionality, rendering single-agent RL approaches intractable.
-As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each driver is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives.
-MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (Jin et al.
-2025).
-Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to single-agent RL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents.
-However, DTDE methods rely on the independent assumption that each agent acts based only on its local information and treats other agents as part of the environment.
-This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (Hu, Feng, and Li 2025), ultimately resulting in suboptimal and inefficient learning.
-To address this problem, we propose two multi-agent policy optimization methods that do not require estimating Q-values or V-values.
-First, we adapt Group Relative Policy Optimization (GRPO) (Shao et al.
-2024) for the ride-sharing task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (Schulman et al.
-2017) with the group average reward.
+Although order dispatch is fundamentally a centralized task—requiring the platform to aggregate information from all vehicles and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (sivagnanam2024multi)) pose a serious Curse of Dimensionality (CoD), rendering Single-Agent Reinforcement Learning (SARL) approaches intractable.
+As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each vehicle is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives.
+MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (jin2025comprehensive).
+Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to SARL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents.
+However, DTDE methods rely on the independent assumption that each agent acts based only on its local observation and treats other agents as part of the environment.
+This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (hu2025bmg), ultimately resulting in suboptimal and inefficient learning.
+To address this problem, we propose two novel multi-agent policy optimization methods that leverage the homogeneous characteristics of agents in the ride-sharing system with AVs, eliminating the need to estimate Q-values or V-values.
+First, we adapt Group Relative Policy Optimization (GRPO) (GRPO) for the order dispatch task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (PPO) with the group average reward-to-go.
 This modification significantly reduces the computational resources required for training and minimizes bias by eliminating neural network estimation errors associated with V-values and Q-values.
-Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step rewards by leveraging the homogeneous characteristics of agents in the ride-sharing task.
-Both methods are validated through case studies using real-world ride-hailing data and demonstrate superior performance compared to state-of-the-art approaches.
+Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step group rewards.
+Both methods are validated through case studies using real-world ride-hailing data and demonstrate promising performance with high efficiency.
 The main contributions of this paper are summarized as follows:
-  • We introduce a novel ride-sharing framework based on GRPO.
-By replacing the baseline in PPO with group average reward, our method eliminates the need to estimate V-values.
+  • We introduce a novel order dispatch method for ride-sharing system with AVs based on GRPO.
+By replacing the baseline in PPO with group average reward-to-go, our method eliminates the need to estimate V-values.
 It successfully improves training performance and stability by removing the influence of estimation bias.
-To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing platforms.
-  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in ride sharing task, the advantage function can be replaced by the group relative reward within only one step.
-To the best of our knowledge, it is the simplest and most efficient RL method for the ride-sharing task.
+To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing systems.
+  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in AV-based ride-sharing system, the advantage function can be replaced by the group relative reward within only one step.
+To the best of our knowledge, it is the simplest and most efficient RL method for the order dispatch task.
 Furthermore, we also specify the conditions under which OSPO can be applied in identical MARL scenarios.
@@
 Furthermore, OSPO surpasses GRPO in performance, which can be attributed to its elimination of the bias introduced by the bounded simulation time present in GRPO.
-Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each driver-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using PPO-style policy gradients.
-Preliminary
-Group Relative Policy Optimization (GRPO)
+2 Preliminary
+2.1 Problem Setup
+In this paper, we consider the optimal order assignment task of an ride-sharing platform with $n$ homogeneous AVs (agents), each equipped with capacity $c$.
+At each time step $t$, there will be $m_{t}$ newly arrived orders from passengers in different locations.
+The platform must dynamically determine how to assign each order to a suitable vehicle in real-time.
+The dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users.
+Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (littman1994markov), defined as $n,S,U,\mathcal{T},R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor.
+The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e.
+information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$.
+In our order dispatch task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and estimated remaining delivery time.
+The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$.
+At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$.
+The transition probability function $\mathcal{T}$ is not modeled explicitly in model-free RL methods.
+At last, the reward function is shaped following (hu2025bmg):
+[…]
+where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the vehicle operational costs, respectively.
+The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively.
+This reward function is designed to comprehensively consider the interests of the platform and passengers, mimicking the operation of a real-world ride-sharing platform.
+It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively.
+When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route.
+Since order dispatch is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
+2.2 Homogeneous Property in Ride-Sharing System
+In this paper, we leverage the homogeneous property of agents (AVs), which is foundational to our proposed methods.
+This homogeneity can be understood from both the task and the MARL perspectives.
+From the task perspective, within the order dispatch system, the primary differences between agents are their spatial locations and the assigned en-route orders.
+The platform itself is agnostic to individual agent identity.
+From the MARL perspective, all agents share an identical structure for their observation space, action space, and reward function.
+Furthermore, the state transition dynamics are governed by the same environmental rules.
+This makes the order dispatch task a fully cooperative one for the platform, where agents can share a common policy.
+This homogeneous property is also a key enabler for parameter and experience sharing, a common practice in prior works (al2019deeppool; enders2023hybrid; hu2025bmg).
+Considering a multi-agent system, the state value function among agents exhibits the following relationship:
+$V_{\pi}(s_{1,t})+\epsilon_{1,t}=V_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (2)
+$=V_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[V_{\pi}(s_{i,t})]\ ,$
+where $\epsilon_{i,t}$ is a residual term, defined as $\epsilon_{i,t}:=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[V_{\pi}(s_{i,t})]-V_{\pi}(s_{i,t})$, and $\mathcal{I}=\{1,2,3,\ldots,n\}$.
+Given the homogeneous property among agents in our scenario, the residual term $\epsilon_{i,t}$ is very small.
+This aligns with the principles of Mean Field MARL (yang2018mean), which has also been adapted for order dispatch tasks in ride-sharing systems (li2019efficient).
+2.3 Group Relative Policy Optimization (GRPO)
 With the success and popularity of Reinforcement Learning from Human Feedback (RLHF), RL-based LLM post-training has garnered increasing attention.
-This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (Rafailov et al.
-2023), Reinforce Leave-One-Out (RLOO) (Ahmadian et al.
-2024), and GRPO (Shao et al.
-2024).
-Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (Shao et al.
-2024) and DeepSeek-V3 (Liu et al.
-2024).
-Serving as a PG (Sutton et al. 1999) based method, the optimization objective of GRPO has a similar format to PPO(Schulman et al. 2017), expressed as:
-$\text{J}_{GRPO}(\theta)=\mathbb{E}_{s,u\sim\pi_{\theta^{-}}}\left[\min\left\{\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)}\hat{\text{A}}_{\pi_{\theta}}(s,u),\text{CLIP}\right.\right.$ (1)
-$\left.\left(\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg{]}-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
-where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (Yu et al.
-2025), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
+This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (rafailov2023direct), Reinforce Leave-One-Out (RLOO) (ahmadian2024back), and GRPO (GRPO).
+Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (GRPO) and DeepSeek-V3 (liu2024deepseek).
+Serving as a PG (sutton1999policy) based method, the optimization objective of GRPO has a similar format to PPO (PPO), expressed as:
+[…]
+$\left.\left(\frac{\pi_{\theta}(u|s)}{\pi_{\theta^{-}}(u|s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg]-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
+where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (yu2025dapo), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
 Aside from the KL divergence term, the core difference between GRPO and PPO lies in the definition of the advantage function. In PPO, the original advantage function is defined as:
@@
 […]
-Problem Setup for Ride-sharing Order Dispatch
-In this paper, we consider the optimal order assignment task of an on-demand ride-sharing platform with $n$ homogeneous drivers (agents), each equipped with a vehicle of capacity $c$.
-At each time step $t$, there will be $m_{t}$ newly arrived orders from customers in different locations.
-The platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suit- able vehicle.
-These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users.
-Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (Littman 1994), defined as $<n,S,U,P,R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor.
-The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e.
-information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$.
-In our ride-sharing task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and remaining delivery time.
-The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$.
-At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$.
-The transition probability function $P$ is not modeled explicitly in model-free RL methods.
-For the reward function $\text{r}(s_{i},u_{i})$, we leave the detailed design in appendix.
-Since ride-sharing is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
-Methodology
-In this section, we first describe how we adapt GRPO for the multi-agent ride-sharing scenario, and analyze the theoretical foundations that support its applicability.
-We then introduce our proposed OSPO method and specify the conditions under which it can be effectively applied.
-GRPO for Ride Sharing
-The workflow of GRPO for ride sharing is shown in Fig.
-1, following a design similar to previous SAC-based work (Enders et al.
-2023).
-Firstly, we require the policy network $\pi_{\theta}$ to directly output the action probability of each agent, expressed as:
+3 Methodology
+In this section, we first introduce how to adapt policy-based MARL methods for the order assignment task in the ride-sharing system (operation stage).
+Then, we propose our GRPO and OSPO methods as optimization solutions (training stage), accompanied by detailed theoretical analysis.
+3.1 Policy-Based Method for Order Dispatch in Ride-Sharing System
+In policy-based deep RL, at each time step $t$, the agent directly generates the probabilities of each action based on the observation:
 $p_{i,t}=\pi_{\theta}([s_{i,t},s^{o}_{t}])\ ,$ (6)
@@
 Due to the homogeneity of the agents, we utilize a shared network among all agents, similar to approaches taken in previous works.
-Considering the action constraint that an order cannot be assigned to multiple agents, we cannot simply choose the action independently for each agent.
-Instead, we view the action probabilities as scores and use bipartite matching to maximize the assignment score, following (Enders et al.
-2023; Hu, Feng, and Li 2025):
+However, in the order dispatch task, we cannot assign the same order to multiple agents, making it infeasible to directly sample actions as $u_{i,t}\sim p_{i,t}$.
+To address this problem, we follow previous policy-based ride-sharing methods (enders2023hybrid; hoppe2024global) by viewing the probabilities as scores and using bipartite matching to maximize the assignment score:
 $\max_{U_{t}}\sum_{i\in\mathcal{I}}u_{i,j,t}\cdot y_{i,j,t},$ (7a)
@@
 $u_{i,j,t}\in\{0,1\},\quad\forall i\in\mathcal{I},j\in\mathcal{J}_{t},$ (7d)
-where $u_{i,j,t}$ indicates whether agent $i$ is assigned order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=-\infty$ for all unavailable workers at time $t$).
-The set =$\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned.
+where $u_{i,j,t}$ indicates whether agent $i$ is assigned to order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=$-$\infty$ for all unavailable workers at time $t$).
+The set $\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned.
 Constraint (7b) ensures that an order can be assigned to at most one agent, while constraint (7c) guarantees that each agent is assigned at most one order.
-After that, we can obtain the reward according to Eq.
-13 and then optimize the policy by maximizing Eq.
+This process is illustrated in the first two steps of Fig.
 1.
-To make GRPO work in our homogeneous cooperative MARL scenario, only two minor changes should be made to the original objective function in the SARL scenario.
-First, when calculating the advantage function, we use the rewards of all agents over all steps to compute the mean value $\mu$ and standard deviation $\sigma$ in Eq.
-3.
-(For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$.
-This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.) Second, for the KL normalization term in Eq.
-1, since we do not have a reference policy like the LLM post-training, we propose to regularly track policy performance and retain the best checkpoint.
-This best policy serves as the reference policy, helping to ensure that the current policy does not deviate significantly from the best one.
-A detailed algorithm is provided in appendix.
-Next, we analyze why our modification to GRPO is valid, allowing the transfer from SARL to MARL. For the advantage function (Eq. 3) in GRPO, we can express it as:
-$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$ (6)
-For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (Schulman et al.
-2017).
-For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 00 as the number of sampled trajectories increases (i.e., as the group becomes larger).
-The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (Wang et al.
-2023).
+3.2 GRPO Solution
+To motivate the adaptation of GRPO to our order dispatch task, we first analyze its theoretical underpinnings. For the advantage function (Eq. 5) in GRPO, we can express it as:
+$\hat{\text{A}}_{\pi}(s_{t},u_{t})$ $=\sum_{\tau\in K_{t}}\gamma^{\tau}\frac{\text{R}(s_{t+\tau},u_{t+\tau})-\mu}{\sigma}$ (8)
+$=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$
+For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (PPO).
+For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 0 as the number of sampled trajectories increases (i.e., as the group becomes larger).
+The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (wang2023math).
 An important feature of such a reward function is that each step reward has a similar range, leading to: $\mathbb{E}_{s_{t},u_{t}\sim\pi}[\text{R}(s_{t},u_{t})]=\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi_{t}$, where $\xi_{t}$ is a small residual and $\mathbb{E}_{t}[\xi_{t}]=0$.
 Thus, we have:
-$\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (7)
+$\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})=\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (9)
 $=\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s_{\tau},u_{\tau}\sim\pi}[\text{R}(s_{\tau},u_{\tau})]-\xi_{\tau}+\xi)$
@@
 Compared to the advantage in PPO (Eq.
-2), there is only one additional factor $\frac{1}{\sigma}$.
-According to (Kurin et al.
-2022), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training.
-Furthermore, based on the theory of policy gradients (Sutton et al.
-1999), adding any baseline or positive factor does not influence the policy direction, regardless of the action.
+4), there is only one additional factor $\frac{1}{\sigma}$.
+According to (kurin2022defense), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training.
 As a result, we consider the advantage of GRPO (Eq.
-3) to be a valid replacement for the advantage in PPO (Eq.
-2).
-One Step Policy Optimization (OSPO)
-We note that in the ride-sharability platform, all vehicles have identical capacity and are assigned to customers purely based on spatial proximity, without any additional attributes or preferences influencing the dispatch decisions.
-As a result, in the long run, the cumulative contribution or average earnings of each driver per unit time must converge to the same value.
-This is because the inherent randomness in order arrivals and assignments is experienced equally by all drivers over time.
-Therefore, it is reasonable to expect that different agents will exhibit similar value functions at each time step, regardless of their individual states.
-In this case, we have the following property:
-$\text{V}_{\pi}(s_{1,t})+\epsilon_{1,t}=\text{V}_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (8)
-$=\text{V}_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[\text{V}_{\pi}(s_{i,t})]\ ,$
-where $\epsilon_{i,t}$ is a small residual term, with $\mathbb{E}_{i\in\mathcal{I}}[\epsilon_{i,t}]=0$.
-As a result, when we ignore the residual terms, the different agents do not differ significantly from the group in GRPO, making it feasible in our scenario.
-Based on the property in Eq.
-8, we can recalculate the advantage function as follows:
+5) to be a valid replacement for the advantage in PPO (Eq.
+4).
+In MARL, the concept of agents is analogous to the group in GRPO, prompting us to consider whether we can adapt GRPO to the MARL scenario, given its high efficiency and promising performance.
+The answer is yes, if the agents are completely homogeneous in a cooperative MARL context, such as in our order dispatch task within the ride-sharing system.
+Under such circumstances, the advantage function can be modified to:
+$\hat{\text{A}}_{\pi}^{\prime}(s_{i,t},u_{i,t})=\frac{1}{\sigma_{t:}}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})-\mu_{t:}\right)\ ,$ (10)
+$\mu_{t:}=\mathbb{E}_{i\in\mathcal{I}}\left[\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})\right]\ .$
+$\sigma_{t:}=\sqrt{\mathbb{E}_{i\in\mathcal{I}}\left[(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})-\mu_{t:})^{2}\right]}\ .$
+In a general MARL scenario, we do not have the property that each step’s reward has a similar range, making it infeasible to utilize $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$ to estimate the V-value as shown in Eq.
+9.
+However, based on the homogeneous property outlined in Eq.
+2, we can use the expectation of the reward-to-go for all agents, $\mu_{t:}$, as an approximation of the V-value.
+Finally, for the KL regularization term in the GRPO objective (Eq.
+3), which originally uses a reference LLM policy $\pi_{ref}$, we innovatively propose to use the best-performing policy checkpoint obtained during training.
+This serves as a stabilizing anchor, preventing the policy from deviating drastically from known good behavior.
+3.3 One Step Policy Optimization (OSPO) Solution
+Based on the homogeneous property in Eq. 2, we can recalculate the advantage function as follows:
 $\text{A}_{\pi}(s_{i,t},u_{i,t})=\text{Q}_{\pi}(s_{i,t},u_{i,t})-\text{V}_{\pi}(s_{i,t})$ (11)
@@
 $=$ $\text{r}(s_{i,t},u_{i,t})-\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}\ .$
-If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (Kurin et al.
-2022), and ignoring the residual term, we arrive at the advantage function:
+If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (kurin2022defense), and ignoring the residual term, we arrive at the advantage function:
 $\hat{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu}{\sigma}\ .$ (12)
 This indicates that we can use only the one-step reward for policy optimization, which we term One Step Policy Optimization (OSPO).
-Even though the format may seem counter-intuitive, we provide an intuitive explanation.
-The validity of Eq.
-10 is based on Eq.
-8: the V-values remain similar among agents at each step due to the homogeneity in the ride-sharing system.
-For any arbitrary agent, regardless of its actions in the current step, it can achieve a similar V-value as others in the next step.
-Therefore, under the independent assumption, the goal of each agent is to maximize its single-step reward.
-In Eq.
-8, removing the baseline and factor leaves only the single-step reward $\text{r}(s_{i,t},u_{i,t})$, which aligns with our intuition.
-Compared to GRPO, our OSPO further improves computational efficiency.
-Moreover, our method demonstrates greater robustness since it is not influenced by the discount factor.
+It is important to note that our method fundamentally differs from myopic methods, even though the advantage function incorporates only one-step rewards.
+The myopic method focuses solely on maximizing the global reward for the current step, which is not the objective of OSPO.
+In contrast, OSPO optimizes the policy of all agents simultaneously by increasing the probability of favorable actions while reducing the probability of unfavorable ones.
+The goal of OSPO remains to maximize the target function defined in Eq.
+3 (a variant of the long-term accumulative reward in standard RL), but it does so in a one-step format that leverages the homogeneous properties.
 Additionally, we propose two changes to the advantage function. First, we suggest directly using the mean value and standard deviation at each step $t$ to replace the values over the entire episode:
-$\bar{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (11)
+$\tilde{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (13)
 This approach not only makes $\mu_{t}$ closer to $\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]$ in Eq.
-9, but also further improves computational efficiency, achieving true one-step computation.
+11, but also further improves computational efficiency, achieving true one-step computation.
+(For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$.
+This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.)
 Second, during training, to better align with the property described in Eq.
-8, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$.
+2, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$.
 This penalty is designed to reduce the utility gap (reflecting the V-value) between agents.
 Consequently, the final advantage function is expressed as:
-$\tilde{\text{A}}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (12)
-where $\alpha$ is a hyper-parameter. (Note that when Eq. 8 is satisfied, $\Delta\delta_{t}$ is close to 0. We also include the term $-\alpha\Delta\delta_{t}$ in the advantage function of GRPO.)
-Experiment
-Experiment Setup
-To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (Taxi and Commission 2024).
+$\tilde{\text{A}}_{\pi}^{\prime}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (14)
+where $\alpha$ is a hyper-parameter.
+Note that when Eq.
+2 is satisfied, $\Delta\delta_{t}$ is close to 0.
+(We also include the term $\alpha\Delta\delta_{t}$ in the advantage function of GRPO to enhance the homogeneous property.) A detailed algorithm process is provided in Appendix A.
+Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each vehicle-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using policy gradients.
+4 Experiment
+4.1 Experiment Setup
+To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (TLCData).
 The data collected from 19:00 to 19:30 on July 17, 2024, is used as the training set, which includes 3,726 valid orders.
-The data from other periods of that day is used as the testing set.
-For the simulation, we set the total number of drivers to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h.
-The optimal routing is achieved using the OSRM simulator (Luxen and Vetter 2011).
-The experiments are conducted using the PyTorch framework (Paszke et al.
-2019) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card.
-Due to device limitations, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode.
+We then evaluate the trained model using data from the entire day of July 18, 2024, selecting the first 30 minutes of data from each hour, resulting in a total of 24 independent tests.
+An expanded generalization experiment is provided in Appendix B.3, using the data from Queens, New York.
+For the simulation, we set the total number of vehicles to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h.
+The optimal routing is achieved using the OSRM simulator (luxen-vetter-2011).
+The experiments are conducted using the PyTorch framework (paszke2019pytorch) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card.
+Following the setup of pervious works, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode.
 We maximize the training episodes at 1,000 and evaluate model performance every 10 episodes.
-To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (Maas et al.
-2013) as the activation function.
-Considering the variable action space (since the number of orders changes each time), we set the network input as a single driver-order pair, ignoring the relationships between orders.
+To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (maas2013rectifier) as the activation function.
+Considering the variable action space (since the number of orders changes each time), we set the network input as a single vehicle-order pair, ignoring the relationships between orders.
 Finally, a softmax layer is used to normalize the output probabilities.
 During training, our models utilized approximately 3.8GB (OSPO) and 5.2GB (GRPO) of GPU memory, with a batch size of 256.
-For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (Hu, Feng, and Li 2025).
-For optimization, we use the Adam optimizer (Kingma 2014) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
+For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (hu2025bmg; hoppe2024global; enders2023hybrid; al2019deeppool).
+For optimization, we use the Adam optimizer (kingma2014adam) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
 In the following, we first compare our GRPO and OSPO ride-sharing methods with previously popular methods of different types, including DTDE, CTDE, and CTCE.
-Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (De Witt et al.
-2020), Multi-Agent PPO (MAPPO) (Yu et al.
-2022), and Independent PG (IPG) (Sutton et al.
-1999).
-Comparative Experiment
-Figure 2: Training Process: Each method is trained three times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$.
-The shaded area represents the range of fluctuations, while the solid line indicates the average value.
+Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (de2020independent), Multi-Agent PPO (MAPPO) (yu2022surprising), and Independent PG (IPG) (sutton1999policy).
+Table 1: Comparison of Different Ride Sharing Methods: Bold entries represent the best results.
+Method OSPO GRPO DeepPool BMG-Q HIVES Enders et al.
+CEVD Hoppe et al.
+Reward Type Local Local Local Local Global Local Global Mixed RL Algorithm OSPO GRPO IDDQN IDDQN QMIX MASAC VD COMA Network Backbone MLP MLP MLP GAT GRU MLP MLP MLP Model Size 20K 20K 20K 117K 16M 118K 23K 167K GPU Occupation (GB) 3.82 5.17 3.97 4.28 6.01 8.19 21.45 9.78 Reward ($\times 10^{3}$) 11.00 ± 2.63 10.90 ± 2.61 10.41 ± 2.34 10.88 ± 2.35 10.70 ± 2.32 9.77 ± 2.41 10.71 ± 2.36 9.93 ± 2.43 Service Rate 1.00 ± 0.00 1.00 ± 0.00 0.99 ± 0.02 1.00 ± 0.01 0.99 ± 0.02 0.99 ± 0.02 1.00 ± 0.01 0.99 ± 0.02 Delivery Time 12.32 ± 0.71 12.40 ± 0.73 12.20 ± 1.43 10.61 ± 0.70 10.81 ± 0.51 12.65 ± 0.52 11.95 ± 0.79 12.54 ± 0.61 Detour Time 1.26 ± 0.41 1.42 ± 0.52 1.24 ± 0.90 0.59 ± 0.59 0.84 ± 0.79 2.72 ± 0.33 1.12 ± 0.85 2.54 ± 0.43 Pickup Time 6.71 ± 0.64 7.33 ± 0.60 8.67 ± 1.23 7.62 ± 0.23 8.13 ± 0.31 8.05 ± 0.18 8.02 ± 0.88 8.10 ± 0.58 Confirmation Time 0.00 ± 0.00 0.00 ± 0.00 0.01 ± 0.03 0.00 ± 0.02 0.00 ± 0.01 0.06 ± 0.20 0.00 ± 0.01 0.02 ± 0.08
+4.2 Experiment Results
+To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories:
+  • DeepPool (al2019deeppool): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (IQL; van2016deep), in which each agent trains its Q-network separately.
+To align it with our experimental scenario, we replace the CNN used in original paper with an MLP similar to our methods (without the softmax).
+  • BMG-Q (hu2025bmg): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (velickovic2017graph) to capture the relationships among neighboring agents.
+  • HIVES (hao2022hierarchical): Based on QMIX (rashid2020monotonic), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
+  • enders2023hybrid: Based on Multi-Agent Soft Actor Critic (MASAC) (haarnoja2018soft), Enders et al.
+propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
+  • CEVD (bose2023sustainable): CEVD modifies the Value Decomposition (VD) paradigm (sunehag2018value) from CTDE to CTCE.
+Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
+  • hoppe2024global: Building on the work of enders2023hybrid, a method based on COMA (foerster2018counterfactual) is developed that utilizes a mixture of global and local rewards.
+Specifically, the global reward encourages cooperation, while the local reward enhances training efficiency.
+For fairness, we modify the state space and reward function to align closely with those used in this paper.
+The training process and testing results are presented in Fig.
+2 and Table 1, respectively.
+We illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order.
+Detailed explanations of these metrics can be found in Appendix B.1.
+According to the results, we observe that GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, which allows them to serve more orders.
+Notably, OSPO, with the smallest network size and lowest GPU utilization, achieves the best performance across most testing scenarios.
+Please note that the results in Table 1 represent the weighted average values from 24 different episodes, calculated according to order volume, along with the standard deviation.
+More details can be found in Appendix B.2.
+Figure 2: Training Process: Each method is trained five times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$.
+The shaded area represents the standard deviation of each method.
 (For delivery time and detour time, only completed orders are counted, as these metrics are uncertain for unfinished orders.)
-Table 1: Model Performance on Testing Sets: Bold entries represent the best results.
-Method Configuration Order Amount in Testing Sets Method Network GPU (GB) 2,850 3,114 3,577 3,910 DeepPool MLP (20K) 3.97 10,690.95±86.93 12,368.83±117.75 12,569.92±77.15 12,990.07±25.07 BMG-Q GAT (117K) 4.28 10,982.45±187.60 12,637.12±116.89 13,064.66±142.65 13,600.33±92.30 HIVES GRU (16M) 6.01 11,044.05±8.30 12,787.39±40.97 12,576.75±46.25 12,097.87±72.32 Enders et al.
-MLP (118K) 8.19 10,036.89±88.64 12,160.56±39.53 12,296.22±188.86 12,128.95±123.22 CEVD MLP (23K) 21.45 10,842.98±31.36 12,702.10±37.39 13,263.08±136.67 13,609.91±17.34 GRPO MLP (20K) 5.17 11,372.81±31.73 13,015.88±78.18 13,644.95±116.94 12,964.86±99.84 OSPO MLP (20K) 3.82 11,467.87±5.39 13,046.27±15.29 13,771.56±81.43 13,429.67±69.79
-To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories.
-A more detailed overview of previous work is provided in the appendix.
-  • DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019) (DTDE): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (Tan 1993; Van Hasselt, Guez, and Silver 2016) and using a CNN as the Q-network.
-To align it with our experimental scenario, we replace the CNN with an MLP similar to our methods (without the softmax).
-  • BMG-Q (Hu, Feng, and Li 2025) (DTDE): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (Velickovic et al.
-2017) to capture the relationships among neighboring agents.
-  • HIVES (Hao and Varakantham 2022) (CTDE): Based on QMIX (Rashid et al. 2020), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
-  • Enders et al.
-(Enders et al.
-2023) (CTDE): Based on Multi-Agent Soft Actor Critc (MASAC) (Haarnoja et al.
-2018), Enders et al.
-propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
-  • CEVD (Bose et al.
-2023) (CTCE): CEVD modifies the Value Decomposition (VD) paradigm (Sunehag et al.
-2018) from CTDE to CTCE.
-Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
-For fairness, we modify the state space and reward function to align closely with those used in this paper.
-Additionally, we also choose a Greedy method as a baseline, which directly matches the order to the closest driver.
-The training process and testing results are shown in Fig.
-2 and Table 1, respectively.
-In Fig.
-2, we illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order.
-We observe that the GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, allowing them to serve more orders.
-In Table 1, it is noteworthy that the OSPO, which has the smallest network size and lowest GPU utilization, achieves the best performance across three testing scenarios.
-However, we also find that the GRPO and OSPO do not perform as well when the order amount increases to 3,910, which is the only scenario where the order amount exceeds 3,726 in the training set.
-Notably, the two methods that perform relatively well in this scenario are BMG-Q (Hu, Feng, and Li 2025) and CEVD (Bose et al.
-2023).
-The primary distinction of these methods compared to others is their use of neighbor information when calculating Q-values.
-In Fig.
-2, we also observe that they achieve lower detour times, suggesting better cooperation among agents through neighbor information.
-These results imply that incorporating more global or neighbor information may be more beneficial in on-pick scenarios.
-Moving forward, it would be advantageous to explore the model’s performance when the training set includes multiple order amount scenarios.
-Ablation Study
+4.3 Ablation Study
 In the ablation study, we primarily investigate the effects of:
   • Reward Normalization: We compare the effect of using single-step reward normalization ($\mu_{t},\sigma_{t}$ in Eq.
-11) versus using the whole episode reward ($\mu,\sigma$ in Eq.
-10) to study the trade-off between computational efficiency and performance.
-  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 12) to determine whether it helps to maintain property (ii) (Eq. 8).
+13) versus using the whole episode reward ($\mu,\sigma$ in Eq.
+12) to study the trade-off between computational efficiency and performance.
+  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 14) to determine whether it helps to maintain homogeneous property (Eq. 2).
   • KL Regulation Term: We assess the effect of using a KL regulation term ($\text{KL}(\pi_{\theta}\|\pi_{ref})$ in Eq.
-1) to illustrate whether the historical best policy can prevent model divergence.
-Figure 3: Result of Ablation Study
+3) to illustrate whether the historical best policy can prevent model divergence.
+Figure 3: Training Process of Ablation Study
 Since it seems unreasonable to use single-step reward normalization in GRPO, we omit this experiment.
@@
 Additionally, we observe that OSPO demonstrates overall better performance than GRPO.
-This is because, although the V-values among agents are generally similar, they have been maintained over a longer time period.
-In our ride-sharing scenario, the number of served orders may differ slightly among agents at each time step, but this variation can be considered negligible over the long term.
-Since our simulation lasts only 30 minutes, this difference becomes more pronounced when calculating the advantage (i.e., using the group average cumulative reward as a baseline introduces a slight bias).
+This advantage arises from the nature of our ride-sharing scenario, where the number of served orders may vary slightly among agents and episodes.
+For instance, an agent may receive a newly assigned order at the last step of the simulation, resulting in a relatively higher reward for that episode.
+While this variation is negligible over the long term, it becomes more pronounced within the confines of our 30-minute simulation when calculating the advantage based on the average reward in Eq.
+8.
 In contrast, OSPO exhibits greater robustness, as it relies solely on one-step rewards and does not encounter this issue.
-Discussions
-In this section, we analyze the current limitations and future directions of our work.
-  • Application Constraints: As mentioned earlier, our method primarily relies on the property that the V-values of different agents remain similar at any time step (Eq.
-8).
-This limits the application scenarios of OSPO to cooperative homogeneous environments.
-It cannot be used in cooperative scenarios where some agents must sacrifice their own interests to maximize the global reward.
-However, this condition can mostly be met in independent MARL, as each agent’s learning goal is to maximize its own long-term reward.
-In the next step, it is worthwhile to explore a wider range of possible application scenarios for OSPO.
-  • Independent Assumption: One important reason for proposing OSPO is to reduce the estimation error of Q-values or V-values in independent MARL scenarios.
-However, the current training approach of our method still follows independent learning.
-The challenges of poor cooperation in independent MARL are yet to be addressed.
-Future work could explore combining OSPO with other methods, such as communication mechanisms or more powerful networks, similar to the approach taken by BMG-Q.
-  • Fairness: Since OSPO is fundamentally based on the similar V-value property (Eq.
-8), and we incorporate a punishment term (Eq.
-12) to encourage the model to adhere to this property, it is important to explore whether this approach contributes to fairness.
-Recently, assignment fairness and discriminatory scenarios have garnered increasing attention in gig systems (Shi et al.
-2021).
-Conclusion
-In this paper, we propose two novel MARL methods for ride-sharing that do not rely on estimating V-values or Q-values.
+5 Conclusion
+In this paper, we propose two novel MARL methods for the order dispatch task in the ride-sharing system with AVs that do not rely on estimating V-values or Q-values.
 First, we successfully adapt GRPO to the ride-sharing context, achieving promising performance by eliminating the estimation errors associated with the V-value network in PPO.
-Building on the insights from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards.
-To the best of our knowledge, OSPO is the simplest and most efficient reinforcement learning method for order dispatch on ride-sharing platforms to date.
+Building on the insights gained from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards.
+To the best of our knowledge, OSPO is the simplest and most efficient RL method for order dispatch on ride-sharing platforms to date.
 Extensive experiments on real-world ride-sharing data demonstrate that our methods consistently outperform previous approaches in most scenarios, while requiring only a simple MLP network and exhibiting low GPU utilization.
-Appendix
-Reward Function Design
-The reward function is shaped following previous work (Hu, Feng, and Li 2025):
-$\text{r}(s_{i},u_{i})$ =$=\begin{cases}\beta_{1}+\beta_{2}p^{in}_{i}-\beta_{3}p^{out}_{i}-\beta_{4}\chi_{i}-\beta_{5}\rho_{i},|u_{i}|=1\\ 0,|u_{i}|=0\end{cases}$ (13)
-where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the payout to drivers, respectively.
-The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively.
-This reward function is designed to comprehensively consider the interests of the platform, drivers, and passengers, mimicking the operation of a real-world ride-sharing platform.
-It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively.
-When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route.
-Algorithm
-The detailed algorithm of our GRPO/OSPO for ride sharing method is illustrated in Algorithm 1.
-Algorithm 1 GRPO for Ride Sharing
+Impact Statement
+We believe this work raises minimal ethical concerns, as the primary focus is on automated vehicles.
+However, in the future, we would like to explore whether our approach could benefit human drivers in ride-sharing systems.
+The core principle of our method is the homogeneous V-value (long-term profit), which may contribute to improving fairness in these systems.
+Appendix A Algorithm
+The detailed algorithm of our GRPO/OSPO for ride-sharing system is illustrated in Algorithm 1.
+Algorithm 1 GRPO/OSPO for Ride-Sharing System
 0: Policy network $\pi_{\theta}$ with shared parameters $\theta$
@@
 0: Environment simulator, Reward function (Eq. 1)
-1:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
-2:  for episode $=1$ to $M$ do
+1:for episode =1=1 to $M$ do
+2:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
 3:  Reset environment, get initial states $\{s_{i,0}\}_{i=1}^{n}$, $s^{o}_{0}$
@@
 9:   Execute assignment $U_{t}$, observe rewards $\{r_{i,t+1}\}_{i=1}^{n}$
-10:        Store transition ($(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
+10:   Store transition $(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
 11:  end for
-12:     Policy Optimization:
-13:     for epoch $=1$ to $M$ do
-14:        Sample batch $\mathcal{B}\sim\mathcal{D}$
-15:        Compute group advantage $\hat{A}_{t}$
-16:        Update policy parameters $\theta$ by maximizing Eq. 1
-17:     end for
-18:     if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
-19:        Update best checkpoint: $\phi\leftarrow\theta$
-20:     end if
-21:  end for
-Related Work
+12:  for epoch =1=1 to $H$ do
+13:   Sample batch $\mathcal{B}\sim\mathcal{D}$
+14:   Compute group advantage $\mathcal{A}_{t}$
+15:   Update policy parameters $\theta$ by maximizing Eq. 3
+16:  end for
+17:  if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
+18:   Update best checkpoint: $\phi\leftarrow\theta$
+19:  end if
+20:end for
+Appendix B Experiment Details
+B.1 Evaluation Metrics
+The evaluation metrics used in this paper include:
+  • Served Rate: The proportion of confirmed trips relative to the total number of trip requests initiated by customers.
+  • Delivery Time: The total time taken to serve a trip from origin to destination.
+  • Detour Time: The additional time spent on delivery beyond the minimum delivery time, which is defined as the time required if the vehicle serves only this trip without bundling others.
+  • Pickup Time: The waiting time for customers between trip confirmation and the arrival of the vehicle at the trip origin.
+  • Confirmation Time: The waiting time for customers from the moment they initiate a trip request to when the platform assigns the trip to a vehicle.
+B.2 Detailed Experiment Results
+Figure 4: Order Temporal Distribution
+Figure 5: Experiment Results in Testing Set
+In Fig.
+4 and 5, we illustrated the order amount and the metrics of different methods under each hour in the whole day.
+First, it’s intuitive that the trend of reward is similar to the order volume, since more orders can bring more profits to platform.
+Then we notice before 8 a.m., when the order volume is very low, the overall performance of different methods is similar, since the idle vehicles are sufficient.
+However, after that, the BMG-Q and HIVES first illustrate slightly advantage during 8 a.m.
+to 1 p.m.
+and then are outperformed by our GRPO and OSPO methods as the order volume further increases after 1 p.m..
+Overall, our GSPO and OSPO mainly outperform others by optimizing the pickup time and confirmation time, leading to the highest order service rate in the on-peak hour 6 p.m..
+Additionally, to further prove our homogeneity assumption, we illustrate the relationship among initial state (location) and metrics regarding cumulative reward (without discount factor $\gamma$) and reward-to-go (with discount factor $\gamma$) of our method.
+The results suggest that in our method, the state does not significantly influence the long-term cumulative reward and reward-to-go.
+As shown in Fig.
+6, this supports our homogeneity assumption.
+Figure 6: The relationship between initial location and metrics about cumulative reward and reward-to-go.
+B.3 Expanded Experiment
+Table 2: Performance of Different Methods in Queens, New York City (TLCData)
+Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool 5222.85 0.64 11.24 1.84 12.30 0.21 BMG-Q 5362.00 0.66 9.63 1.08 12.98 0.27 HIVES 3560.80 0.60 8.30 0.36 14.67 0.41 Enders et al.
+4543.68 0.61 10.41 0.85 13.39 2.25 CEVD 4388.83 0.62 11.61 1.33 13.74 0.29 Hoppe et al.
+4728.93 0.63 9.78 0.53 13.45 2.18 GRPO 5365.89 0.64 10.56 1.86 12.21 0.12 OSPO 5521.74 0.67 11.37 2.21 11.88 0.21
+To further demonstrate the generalization of our method, we conduct experiments using High Volume For-Hire Vehicle (FHV) trip data from Queens, New York City (TLCData).
+Unlike the Manhattan data used in our previous experiment, Queens is a larger area with a lower order density.
+This difference in order distribution can complicate the process of determining which orders to bundle together for vehicle sharing.
+In this scenario, we select the data from 19:00 to 19:30 on July 17, 2024, which includes 2,024 valid orders and 500 vehicles.
+The detailed experimental results are presented in Table 2.
+The results demonstrate that our proposed GRPO and OSPO methods continue to exhibit superior performance, primarily by optimizing pickup and confirmation times.
+Specifically, OSPO significantly outperforms the other methods, including GRPO.
+This advantage arises because, in a more challenging and dynamic scenario (due to the enlarged area), the GRPO method is more severely impacted by errors associated with the bounded time horizon, whereas OSPO leverages a one-step reward to mitigate this shortcoming.
+Appendix C Related Work
 In this section, we provide a brief overview of current ride-sharing methods.
-Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019), independent DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods.
+Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (al2019deeppool), DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods.
 Subsequent works have primarily aimed to enhance algorithms by fostering better cooperation among agents.
-For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (Hu, Feng, and Li 2025).
-Li et al.
-(Li et al.
-2019) also propose using mean field approaches to capture more useful global information.
+For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (hu2025bmg).
+li2019efficient also propose using mean field approaches to capture more useful global information.
 Additionally, some works attempt to transfer the successes of CTDE and CTCE to ride-sharing, where centralized critics can facilitate better cooperation among agents.
 However, these approaches face various shortcomings and challenges.
-Lima et al.
-(De Lima et al.
-2020) first introduce QMIX, but only consider tasks involving a limited number of drivers in simulated grid scenarios, which presents a significant gap to real-world applications.
-Hao et al.
-(Hao and Varakantham 2022) explore hierarchical structures to address the curse of dimensionality in QMIX (Rashid et al.
-2020), but the input dimensions of the mixture network still increase with the number of agents.
-Bose et al.
-(Bose et al.
-2023) propose a Value Decomposition (VD) (Sunehag et al.
-2018) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
-Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (Zhang et al.
-2024), order bundling (Jiang, Xu, and Varakantham 2025), and multi-modal transportation (Hu, Dong, and Li 2025).
+de2020efficient first introduce QMIX to ride-sharing, but only consider tasks involving a limited number of vehicles in simulated grid scenarios, which presents a significant gap to real-world applications.
+hao2022hierarchical explore hierarchical structures to address the curse of dimensionality in QMIX (rashid2020monotonic), but the input dimensions of the mixture network still increase with the number of agents.
+bose2023sustainable propose a Value Decomposition (VD) (sunehag2018value) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
+Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (zhang2024joint), order bundling (jiang2025optimizing), and multi-modal transportation (hu2025coordinating).
 Since these topics extend beyond the scope of this paper, we do not discuss them further here.
-Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (Hu, Feng, and Li 2025).
+Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (hu2025bmg).
 Our work addresses this gap by proposing GRPO/OSPO—eliminating value estimation entirely through group rewards—for stable and scalable coordination in homogeneous MARL systems.